# A Vision-Language Pre-training Model-Guided Approach for Mitigating Backdoor Attacks in Federated Learning

### v1

@@ Introduction
Clustering presents a potential solution for distinguishing between malicious and benign models in FL. In CLIP-Fed, we adopt HDBSCAN for pre-aggregation filtering, following advanced methods (Nguyen et al. 2022). Because high-dimensional parameter spaces suffer from the curse of dimensionality, which reduces the reliability of distance metrics like cosine similarity, we use PCA to project model parameters into a lower-dimensional space where distance computations are more meaningful and robust. Following the threat assumptions and parameter settings established in Flame (Nguyen et al. 2022) and FreqFed (Fereidooni et al. 2024), we operate under the premise that benign models constitute the majority of clients while malicious models remain in the minority. Accordingly, we set the minimum cluster size to encompass at least 50% of clients, ensuring the resulting clusters contain predominantly benign model updates. While this clustering process enables CLIP-Fed to eliminate most malicious models prior to aggregation, it is acknowledged that the identified benign clusters may still contain some residual malicious models. After model clustering, CLIP-Fed performs global model aggregation as specified in:
$\theta_{s}^{t+1}=\sum_{i\in S_{t}}\frac{D_{i}}{D_{S_{t}}}\theta_{i}^{t},\ \theta\in\{f,c\},$ (2)
@@
Table 1: Performance comparison between CLIP-Fed and other defense methods.
Feature Rectification via Prototype Alignment
In this phase, CLIP-Fed utilizes the CLIP model, which is equipped with rich prior knowledge, as an auxiliary module to guide the rectification and retraining of the feature extraction module of global model on the server. In backdoor models, the feature extraction module exhibits the increased sensitivity to trigger-associated patterns. Existing research (Shi et al. 2024) has shown that leveraging CLIP’s strongly supervised feature prototypes can effectively guide the generation of federated features. Inspired by this insight, we employ contrastive learning between CLIP-derived class prototypes and features generated by the global model feature extractor, thus mitigating prototype deviations caused by backdoor trigger interference. As shown in Eq. 3, in the first part, we treat the output of the feature extractor $v_{i}^{c}$ (Sample $i$ belonging to class $c$) and the output of CLIP’s visual encoder $e_{c,\text{clip}}^{I}$ as positive pairs, while considering the output of the feature extractor $v_{i}^{c}$ and other class prototypes $e_{k,\text{extractor}}$ from the feature extractor as negative pairs.
$\mathcal{L}_{i}^{\text{proto-1}}=-\log\frac{\exp\left(sim\langle v_{i}^{c},e_{c,\text{clip}}^{I}\rangle/\tau\right)}{\sum_{k=1[k\neq c]}^{C}\exp\left(sim\langle v_{i}^{c},e_{k,\text{extractor}}\rangle/\tau\right)},$ (3)
@@
$\mathcal{L}_{i}^{\text{proto-2}}=-\log\frac{\exp\left(sim\langle v_{i}^{c},e_{c,\text{clip}}^{T}\rangle/\tau\right)}{\sum_{k=1[k\neq c]}^{C}\exp\left(sim\langle v_{i}^{c},e_{k,\text{clip}}^{T}\rangle/\tau\right)}.$ (4)
The total prototype contrastive loss as shown in:
$\mathcal{L}_{\text{pcl}}=\sum_{i=1}^{C\times m}\left(\mathcal{L}_{i}^{\text{proto-1}}+\mathcal{L}_{i}^{\text{proto-2}}\right).$ (5)
The proposed approach systematically eliminates backdoor-induced deviations in feature representations while maintaining the model’s core classification performance.
Global Model Knowledge Transfer
To thoroughly eliminate residual backdoor patterns in the global model, CLIP-Fed employs knowledge transfer to purify the global model using CLIP’s prior knowledge. In this framework, CLIP acts as a “teacher” model, while the global model serves as the “student” model. During each iteration, CLIP provides guidance to the global model using the server dataset, progressively reducing the impact of the backdoor. We utilize the KL divergence to measure the discrepancy between logits distributions of the global model and CLIP on the annotated dataset. Specifically, the loss function for the knowledge transfer process is formulated as
$\mathcal{L}_{\text{kt}}=\mathcal{L}_{ce}\left(y,l_{\text{global}}\right)+\beta\cdot\mathcal{D}_{\text{KL}}\left(l_{\text{clip}}\|l_{\text{global}}\right),$ (6)
@@
Experiments
@@
Table 2: Performance comparison between CLIP-Fed w/o FR, CLIP-Fed w/o KD, and CLIP-Fed.
@@
Figure 3: The visualization of the feature representations under different attacks on CIFAR-10-LT with/without CLIP-Fed.
@@
Figure 4: Comparison of MA and ASR against BadNets attacks with varying malicious client proportions under CIFAR-10-LT.
@@
Table 3: Performance comparison against BadNets attacks with different proportions of malicious clients under CIFAR-10-LT.
@@
Table 4: Performance of CLIP-Fed against BadNets attacks with different degrees of Non-IID.
@@
Table 5: Defense performance against BadNets attacks with different attack strength under CIFAR-10-LT.
Datasets. We evaluated CLIP-Fed on CIFAR-10, CIFAR-10-LT, and CIFAR-100 (Krizhevsky 2009). CIFAR-10-LT introduced an exponential class imbalance to simulate real-world long-tailed distributions. To emulate Non-IID client settings, we adopted Pathological Non-IID partitioning (Blanchard et al. 2017).
@@
Implementation Details. We simulated a FL scenario with 100 clients, where a portion is malicious. Malicious clients conducted four types of backdoor attacks, including BadNets (Gu, Dolan-Gavitt, and Garg 2017), DBA (Xie et al. 2020), LayerAttack, and AFA, following an “all-to-one” strategy. ResNet-18 was used as the model backbone. All experiments were conducted on an RTX 4090 GPU with Python 3.9, PyTorch 2.2.2, and CUDA 11.8. Unless stated otherwise, we simulated Non-IID client data distributions and employed the ViT-L/14 version of CLIP.
@@
Baselines and Existing Defenses. We compared CLIP-Fed with representative backdoor defense methods in FL, including: (1) FedAvg (McMahan et al. 2017), a standard aggregation baseline without defense, (2) FLAME (Nguyen et al. 2022) that combines clustering and differential privacy, (3) FLTrust (Cao et al. 2021) that is a server-side validation-based defense, (4) FoolsGold (Fung, Yoon, and Beschastnikh 2020), a defense scheme based on the similarity of contributions among clients, and (5) FEDCPA (Han et al. 2023) that focuses on critical parameter similarity.
Performance Comparison
@@
Table 1 presented a comparison of Main Task Accuracy (MA) and Attack Success Rate (ASR) across different methods under varying data distributions and backdoor attack. The results demonstrated that CLIP-Fed consistently achieved effective mitigation of backdoor across various attacks and datasets. Specifically, CLIP-Fed achieved the lowest ASR under all four attack types on both the CIFAR-10 and CIFAR-10-LT datasets. On the CIFAR-10, our scheme achieved ASR reductions of 0.78%, 2.36%, 2.35%, and 2.63% under BadNets, DBA, LayerAttack, and AFA, respectively, compared to the SOTA defense FLAME. On the CIFAR-10-LT dataset that exhibited a significant long-tail distribution, our scheme reduced ASR by 0.15%, 0.09%, 3.53%, and 1.61% under the same four attacks, compared to the SOTA defense FEDCPA. In terms of MA, our scheme also had advantages over existing solutions, as it achieved backdoor defense without significantly compromising task performance. This indicated that CLIP-Fed was able to accurately filter out malicious models while preserving the contributions of benign ones, making it more effective than standalone pre-aggregation detection or post-aggregation noise injection methods. On the CIFAR100 dataset with a higher-level complexity of classification tasks, CLIP-Fed achieved the best defense under the four types of attacks.
Ablation Study
@@
We evaluated two variants of CLIP-Fed: CLIP-Fed w/o FR (without feature extractor rectification), and CLIP-Fed w/o KD (without knowledge distillation). Table 2 compared the accuracy of the global model and the defense performance under identical experiment setups. We observed that both CLIP-Fed w/o FR and CLIP-Fed w/o KD achieved lower ASR than FedAvg (as shown in Table 1) across all three datasets and four types of attacks. The results indicated that individually applying either feature extractor retraining or global model knowledge distillation could partially reduce the association between trigger patterns and target labels, thereby mitigating backdoor effects. Moreover, CLIP-Fed combined both feature retraining and knowledge distillation, such that it achieved the highest MA and lowest ASR across all datasets and attack methods.
Visualization of Backdoor Elimination
@@
As shown in Figure 3, t-SNE was used to visualize the feature space distribution of test samples under BadNets attacks, both with and without CLIP-Fed. We added triggers to 5% of the test samples. In the visualization, samples containing triggers were outlined in black. Figure 3a presented the visualization of test samples after being processed by the CLIP model. The CLIP model demonstrated strong robustness to trigger samples and successfully clustered samples from the same class together in the feature space. This suggested that using CLIP’s prior knowledge to guide FL in backdoor purification was a feasible approach. Figure 3b showed the feature space clustering after a backdoor attack without any defense. We observed that nearly all samples with triggers formed a separate cluster, indicating a successful backdoor attack. Figures 3c and 3d showed the clustering results after training with CLIP-Fed for 50 and 200 epochs, respectively. The trigger samples were no longer isolated and rejoined the clusters of clean samples from the same class. The results indicated that CLIP-Fed disrupted backdoor feature clustering and weakened the link between triggers and target labels, effectively mitigating the backdoor.
Parameter Impact
@@
The Impact of Number of Malicious Clients. Table 3 and Figure 4 presented a comparison of the defense performance of CLIP-Fed and existing methods under varying proportions of malicious clients. When the proportion of malicious clients was 0.2, FEDCPA, FLAME, and CLIP-Fed demonstrated comparable effectiveness in reducing the ASR. However, as the proportion exceeded 0.2, the defense performance of existing methods began to degrade to varying extents. Notably, FEDCPA experienced the most significant drop in performance, with an ASR of 90.22% when the malicious client ratio was 0.3, and 96.54% when it increased to 0.4. In contrast, CLIP-Fed consistently maintained strong defense performance across all settings. When the proportion of malicious clients was 0.3 and 0.4, the ASR of CLIP-Fed remained significantly lower than that of the SOTA method FLAME by 3.47% and 5.66%, respectively. In terms of MA, CLIP-Fed also performed favorably, surpassing FedAvg by 2.53%, 1.75%, and 4.17% at malicious client ratios of 0.2, 0.3, and 0.4, respectively. A single pre-aggregation defense strategy made existing methods vulnerable to high proportions of malicious clients, whereas CLIP-Fed’s post-aggregation backdoor purification supplements the defense capability under such conditions. Overall, compared with existing approaches, CLIP-Fed defended against backdoor attacks across different malicious client ratios while minimizing the degradation of task performance.
@@
The Impact of Non-IID. Table 4 presented the defense performance of CLIP-Fed against BadNets attack under varying degrees of Non-IID data distribution. On the CIFAR-10 dataset, the CLIP-Fed approach achieved an average MA of 80.71% and an average ASR of 1.96%. On the CIFAR-100 dataset, the average MA was 45.67% and the average ASR was 0.51%. Across both datasets, the degree of Non-IID distribution caused only minor fluctuations in MA and ASR, and its impact on defense performance was negligible. This was because heterogeneous client data distributions primarily affected the filtering of malicious models before aggregation, while CLIP-Fed’s backdoor purification strategy was decoupled from model aggregation, thereby eliminating the impact of Non-IID data on purification effectiveness. The above demonstrated that CLIP-Fed was effective under different levels of Non-IID data distribution, maintaining high MA and low ASR for robust backdoor defense.
@@
Defense under Different Attack Strength. Trigger size and poisoning ratio served as crucial factors affecting the attack strength. As shown in Table 5, in the absence of defense, MA gradually declined while ASR increased as the trigger size became larger. Larger triggers resulted in more distinguishable features, making it easier for the model to associate them with the target label. Additionally, higher poisoning rates further intensified the attack, leading to a sharper drop in model performance. When CLIP-Fed was applied, backdoor were effectively prevented from being embedded into the global model, regardless of trigger size or poisoning rate. The connection between trigger features and the target label was severed, and residual backdoor patterns in the global model were removed, making the defense independent of attack strength. The above findings indicated that the influence of trigger size and poisoning rate became minimal when CLIP-Fed was employed.
Conclusion

### v2

@@ 3.3. Dynamic Model Filtering by Clustering
Clustering presents a potential solution for distinguishing between malicious and benign models in FL. In CLIP-Fed, we adopt HDBSCAN for pre-aggregation filtering, following advanced methods (Nguyen et al., 2022). Because high-dimensional parameter spaces suffer from the curse of dimensionality, which reduces the reliability of distance metrics like cosine similarity, we use PCA to project model parameters into a lower-dimensional space where distance computations are more meaningful and robust. Following the threat assumptions and parameter settings established in Flame (Nguyen et al., 2022) and FreqFed (Fereidooni et al., 2024), we operate under the premise that benign models constitute the majority of clients while malicious models remain in the minority. Accordingly, we set the minimum cluster size to encompass at least 50% of clients, ensuring the resulting clusters contain predominantly benign model updates. While this clustering process enables CLIP-Fed to eliminate most malicious models prior to aggregation, it is acknowledged that the identified benign clusters may still contain some residual malicious models. After model clustering, CLIP-Fed performs global model aggregation as specified in Eq. (2):
(2) $\theta_{s}^{t+1}=\sum_{i\in S_{t}}\frac{D_{i}}{D_{S_{t}}}\theta_{i}^{t},\ \theta\in\{f,c\},$
@@
Table 1. Performance comparison between CLIP-Fed and other defense methods
@@
3.4. Feature Rectification via Prototype Alignment
The proposed approach systematically eliminates backdoor-induced deviations in feature representations while maintaining the model’s core classification performance. In general, backdoor attacks involve malicious clients injecting trigger-infused samples with target labels, causing the model to misclassify these inputs. This leads to a shift in the target class prototype, either pulling it toward the attack samples or creating a fake prototype, along with increased intra-class variance. In this phase, CLIP-Fed utilizes the CLIP model, which is equipped with rich prior knowledge, as an auxiliary module to guide the rectification and retraining of the feature extraction module of global model on the server. In backdoor models, the feature extraction module exhibits the increased sensitivity to trigger-associated patterns. Existing research (Shi et al., 2024) has evidenced that using CLIP’s strongly supervised feature prototypes can guide the generation of federated features. Inspired by this fact, we employ contrastive learning between CLIP-derived class prototypes and features generated by the global model feature extractor, thereby mitigating prototype deviations caused by backdoor trigger interference. As shown in Eq. 3, in the first part, we treat the output of the feature extractor $v_{i}^{c}$ (Sample $i$ belonging to class $c$) and the output of CLIP’s visual encoder $e_{c,\text{clip}}^{I}$ as positive pairs, while considering the output of the feature extractor $v_{i}^{c}$ and other class prototypes $e_{k,\text{extractor}}$ from the feature extractor as negative pairs.
(3) $\mathcal{L}_{i}^{\text{proto-1}}=-\log\frac{\exp\left(sim\langle v_{i}^{c},e_{c,\text{clip}}^{I}\rangle/\tau\right)}{\sum_{k=1[k\neq c]}^{C}\exp\left(sim\langle v_{i}^{c},e_{k,\text{extractor}}\rangle/\tau\right)},$
@@ 3.4. Feature Rectification via Prototype Alignment
(4) $\mathcal{L}_{i}^{\text{proto-2}}=-\log\frac{\exp\left(sim\langle v_{i}^{c},e_{c,\text{clip}}^{T}\rangle/\tau\right)}{\sum_{k=1[k\neq c]}^{C}\exp\left(sim\langle v_{i}^{c},e_{k,\text{clip}}^{T}\rangle/\tau\right)}.$
The total prototype contrastive loss as shown in Eq. (5):
(5) $\mathcal{L}_{\text{pcl}}=\sum_{i=1}^{C\times m}\left(\mathcal{L}_{i}^{\text{proto-1}}+\mathcal{L}_{i}^{\text{proto-2}}\right).$
3.5. Global Model Knowledge Transfer
To thoroughly eliminate residual backdoor patterns in the global model, CLIP-Fed employs knowledge transfer to purify the global model by using CLIP’s prior knowledge. CLIP acts as a “teacher” model and the global model serves as the “student” model. During each iteration, CLIP provides guidance to the global model using the server dataset, progressively reducing the impact of the backdoor. We utilize the KL divergence to measure the discrepancy between logits distributions of the global model and CLIP on the annotated dataset. Specifically, the loss function for the knowledge transfer process is formulated as
(6) $\mathcal{L}_{\text{kt}}=\mathcal{L}_{ce}\left(y,l_{\text{global}}\right)+\beta\cdot\mathcal{D}_{\text{KL}}\left(l_{\text{clip}}\|l_{\text{global}}\right),$
@@ 3.5. Global Model Knowledge Transfer
4. Experiments
Table 2. Performance comparison between CLIP-Fed w/o FR, CLIP-Fed w/o KD, and CLIP-Fed
@@ 4. Experiments
Figure 3. The visualization of the feature representations under different attacks on CIFAR-10-LT with/without CLIP-Fed.
(a) Original Image
(b) Image with Trigger
(c) Frequency-perturbed Image
Figure 4. Comparison of original, triggered, and frequency-perturbed images.
Figure 5. Comparison of MSE values between the original image and the image with pixel blocks in different frequency domain intervals.
@@
Figure 6. Comparison of MA and ASR against BadNets attacks with varying malicious client proportions under CIFAR-10-LT.
Table 3. Performance comparison against BadNets attacks with different proportions of malicious clients under CIFAR-10-LT
@@
Table 4. Performance of CLIP-Fed against BadNets attacks with different degrees of Non-IID
@@
Table 5. Defense performance against BadNets attacks with different attack strength under CIFAR-10-LT
@@
Datasets. We evaluated CLIP-Fed on CIFAR-10, CIFAR-10-LT, and CIFAR-100 (Krizhevsky, 2009). CIFAR-10-LT was a variant of CIFAR-10 with long-tail distribution characteristics. The number of its category samples decreased exponentially, thereby simulating the category imbalance phenomenon in the real world. To simulate Non-IID data, we adopted the Pathological Non-IID partitioning (Blanchard et al., 2017).
@@
Implementation Details. We simulated a FL scenario with 100 clients, where a portion is malicious. Malicious clients conducted four types of backdoor attacks, including BadNets (Gu et al., 2017), DBA (Xie et al., 2020), LayerAttack, and AFA, following an “all-to-one” strategy. ResNet-18 was used as the model backbone. All experiments were conducted on an RTX 4090 GPU with Python 3.9, PyTorch 2.2.2, and CUDA 11.8. Unless stated otherwise, we simulated Non-IID client data distributions and employed the ViT-L/14 version of CLIP.
Attack Approach. We evaluated the defense effectiveness of CLIP-Fed against several existing FL backdoor attacks. Specifically, the adversary adopted the following attack strategies. (1) BadNets (Gu et al., 2017), a typical backdoor attack method that directly embedded a static trigger into training samples; (2) Distributed Backdoor Attack (DBA) (Xie et al., 2020), a federated backdoor attack that decomposed a global trigger pattern and embedded its fragments into various clients; (3) LayerAttack, a hierarchical poisoning method targeting critical layers in neural networks, which was particularly effective against distance-based defense mechanisms; (4) AFA, a variant of LayerAttack that performed hierarchical bit flipping on sensitive layers, effective against sign-based defenses. In addition, we assumed that the adversary adopted an “all to one” attack strategy when a backdoor attack was launched, which meant the adversary added triggers to samples of all categories and changed their labels to a target category.
@@
Baselines and Existing Defenses. We compared CLIP-Fed with representative backdoor defense in FL, including FedAvg (McMahan et al., 2017) (a standard aggregation baseline without defense), FLAME (Nguyen et al., 2022) (combining clustering and differential privacy), FLTrust (Cao et al., 2021) (a server-side validation-based defense method), FoolsGold (Fung et al., 2020) (a defense scheme based on the similarity of contributions among clients), and FEDCPA (Han et al., 2023) (focusing on critical parameter similarity).
4.1. Performance Comparison
@@
Table 1 presented a comparison of Main Task Accuracy (MA) and Attack Success Rate (ASR) across various methods under distinct data distributions and backdoor attacks. The results demonstrated that CLIP-Fed consistently achieved effective backdoor mitigation in various attacks and datasets. Specifically, CLIP-Fed had the lowest ASR under all four attack types on both CIFAR-10 and CIFAR-10-LT datasets. On CIFAR-10, our scheme achieved ASR reductions of 0.78%, 2.36%, 2.35%, and 2.63% under BadNets, DBA, LayerAttack, and AFA, respectively, compared to SOTA defense FLAME. On CIFAR-10-LT dataset that had a significant long-tail distribution, our scheme reduced ASR by 0.15%, 0.09%, 3.53%, and 1.61% under the same four attacks, compared to the SOTA defense FEDCPA. In terms of MA, our scheme also had advantages over existing solutions, as it achieved backdoor defense without significantly compromising task performance. This indicated that CLIP-Fed was able to accurately filter out malicious models while preserving contributions of benign ones, performing a higher-level effectiveness than standalone pre-aggregation detection or post-aggregation noise injection methods. On CIFAR100 dataset with a higher-level complexity of classification tasks, CLIP-Fed achieved the best defense under four types of attacks.
4.2. Ablation Study
@@
We conducted comprehensive ablation experiments on CIFAR-10 under backdoor attacks to thoroughly validate the effectiveness of CLIP-Fed’s core components. Two variants of CLIP-Fed were evaluated, which were CLIP-Fed w/o FR (without feature extractor rectification) and CLIP-Fed w/o KD (without knowledge distillation). Table 2 showed comparisons of the accuracy on the global model and the defense performance under identical experiment setups. We observed that both CLIP-Fed w/o FR and CLIP-Fed w/o KD achieved lower ASR than FedAvg (as shown in Table 1) across all three datasets and four types of attacks. The results indicated that individually applying either feature extractor retraining or global model knowledge distillation could partially eliminate the association between trigger patterns and target labels, thereby mitigating backdoor effects. Thus, CLIP-Fed combined both feature retraining and knowledge distillation, so that it achieved the highest MA and lowest ASR in call cases.
4.3. Visualization of Backdoor Elimination
@@
To evaluate the purification effectiveness after multiple rounds of backdoor elimination (feature rectification and global model knowledge transfer), we examined changes in the association between trigger samples and target labels before/after defense. As shown in Figure 3, t-SNE was used to visualize the feature space distribution of test samples under BadNets attacks, both with and without CLIP-Fed. We added triggers to 5% of the test samples. In visualization, samples containing triggers were outlined in a black color. Figure 3a presented the visualization of test samples after being processed by the CLIP model. The CLIP model demonstrated strong robustness to trigger samples and successfully clustered samples from the same class together in the feature space. This suggested that using CLIP’s prior knowledge to guide FL in backdoor purification was a feasible approach. Figure 3b showed the feature space clustering after a backdoor attack without any defense. We observed that most samples with triggers formed a separate cluster, indicating a successful backdoor attack. Figures 3c and 3d showed clustering results after training with CLIP-Fed for 50 and 200 epochs, respectively. The trigger samples were no longer isolated and rejoined the clusters of clean samples from the same class. The results indicated that CLIP-Fed disrupted backdoor feature clustering and weakened the association between triggers and target labels.
4.4. Analysis of the Frequency Sensitivity-guided Image Augmentation
To assess the proposed frequency sensitivity-guided image augmentation, we compared the original image, a pixel-block triggered image, and our frequency-perturbed image, as shown in Figure 4. Visual results demonstrated that our scheme avoided introducing visible distortion, thereby preserving the semantic content. We computed MSE between the DCT spectra of the clean and triggered images across $k$ frequency bands. As shown in Figure 5, the 7th frequency band exhibited the highest MSE, indicating the strongest sensitivity to trigger artifacts. We injected Gaussian noise into $F_{high}(u,v)$, and reconstructed the image via IDCT (refer to Figure 4c). This approach simulated potential backdoor signals without compromising visual quality. Moreover, by focusing perturbations on the most trigger-sensitive frequency regions, the augmented images served as effective surrogates for trigger patterns, enabling the global model to generalize more effectively against diverse backdoor attacks during training.
4.5. Parameter Impact
@@
Impact of Number of Malicious Clients. Table 3 and Figure 6 presented a comparison of the defense performance of CLIP-Fed and existing methods under varying proportions of malicious clients. When the proportion of malicious clients was 0.2, FEDCPA, FLAME, and CLIP-Fed demonstrated comparable effectiveness in reducing ASR. However, as the proportion exceeded 0.2, the defense performance of existing methods began to degrade to varying extents. In particular, FEDCPA experienced the most significant drop in performance, with an ASR of 90.22% when the malicious client ratio was 0.3, and 96.54% when it increased to 0.4. In contrast, CLIP-Fed consistently maintained strong defense performance across all settings. When the proportion of malicious clients was 0.3 and 0.4, the ASR of CLIP-Fed remained significantly lower than that of the SOTA method FLAME by 3.47% and 5.66%, respectively. In terms of MA, CLIP-Fed also performed favorably, surpassing FedAvg by 2.53%, 1.75%, and 4.17% at malicious client ratios of 0.2, 0.3, and 0.4, respectively. As shown in Figure 6, CLIP-Fed consistently demonstrated a clear advantage across all stages of training. A single pre-aggregation defense strategy made existing methods vulnerable to high proportions of malicious clients, whereas CLIP-Fed’s post-aggregation backdoor purification supplements the defense capability under such conditions. Overall, compared with existing approaches, CLIP-Fed effectively defended against backdoor attacks across different malicious client ratios while minimizing the degradation of task performance in FL.
@@
Impact of Non-IID. Table 4 presented the defense performance of CLIP-Fed against BadNets attack under varying degrees of Non-IID data distribution. On CIFAR-10 dataset, the CLIP-Fed approach achieved an average MA of 80.71% and an average ASR of 1.96%. On CIFAR-100 dataset, the average MA was 45.67% and the average ASR was 0.51%. Across both datasets, the degree of Non-IID distribution caused only minor fluctuations in MA and ASR, and its impact on defense performance was negligible. The reason for these phenomena was because heterogeneous client data distributions primarily affected the filtering of malicious models before aggregation, while CLIP-Fed’s backdoor purification strategy was decoupled from model aggregation. The above clearly demonstrated that CLIP-Fed was effective under different levels of Non-IID data distribution, thereby maintaining a high MA and a low ASR for consistent robust backdoor defense.
@@
Defense under Different Attack Strength. To investigate CLIP-Fed’s defensive performance against backdoor attacks with varying strengths, we evaluate attacks employing different trigger sizes and poisoning ratio on the CIFAR-10-LT dataset. Trigger size and poisoning ratio served as crucial factors affecting the attack strength. When attackers became aware of the deployed defense mechanisms, they might avoid defense strategies based on malicious model screening rules, either increasing trigger size and poisoning to enhance attack chance or reducing trigger size and poisoning to improve backdoor concealment. As shown in Table 5, in the absence of defense, MA gradually declined while ASR increased as the trigger size became larger. Larger triggers resulted in more distinguishable features, making it easier for the model to associate them with the target label. Additionally, higher poisoning rates further intensified the attack, resulting in a sharper drop in model performance. When CLIP-Fed was applied, backdoor were effectively prevented from being embedded into the global model, regardless of trigger sizes or poisoning rates. The connection between trigger features and the target label was severed, and residual backdoor patterns in the global model were removed, making the defense independent of attack strength. CLIP-Fed demonstrated consistent effectiveness in mitigating backdoor effects across adaptive attacks with various trigger sizes and poisoning ratios.
Impact of Distinct CLIP Versions on Backdoor Defense Performance. We evaluated the impact of different versions of CLIP (Radford et al., 2021) on the backdoor defense performance. Specifically, we adopted ViT-B/32, ViT-B/16, and ViT-L/14 as VLPs. Figure 7 showed the comparison of MA and ASR on CIFAR-10 dataset under BadNets (Gu et al., 2017) and DBA (Xie et al., 2020) attacks when different CLIP versions were applied. In terms of accuracy, we observed that CLIP-Fed achieved over 80% MA across all CLIP versions, evidencing that the rich prior knowledge from each version was sufficient for the CIFAR-10 dataset. Regarding ASR, we found that the success rate under BadNets attacks was higher than that of DBA attacks, due to the stronger impact of centralized attacks. However, all CLIP versions achieved an ASR below 2% under both types of attacks, demonstrating that they could all effectively support CLIP-Fed in defending against backdoor threats. These findings indicate that the defense performance of CLIP-Fed remains consistently strong across different CLIP versions. This suggested that CLIP-Fed be not highly sensitive to the scale of the VLPs and flexibly adapt to various deployment scenarios without significant degradation in defense effectiveness.
(a) MA with different version of CLIP
(b) ASR with different version of CLIP
Figure 7. Comparison of MA and ASR in resisting BadNets and DBA attacks with different version of CLIP under CIFAR-10.
5. Conclusion